# TraceFL: Interpretability-Driven Debugging in Federated Learning via Neuron Provenance

### v3

@@ Abstract
In Federated Learning, clients train models on local data and send updates to a central server, which aggregates them into a global model using a fusion algorithm. This collaborative yet privacy-preserving training comes at a cost—FL developers face significant challenges in attributing global model predictions to specific clients. Localizing responsible clients is a crucial step towards (a) excluding clients primarily responsible for incorrect predictions and (b) encouraging clients who contributed high-quality models to continue participating in the future. Existing ML explainability approaches are inherently inapplicable as they are designed for single-model, centralized training.
We introduce TraceFL, a fine-grained neuron provenance capturing mechanism that identifies clients responsible for the global model’s prediction by tracking the flow of information from individual clients to the global model. Since inference on different inputs activates a different set of neurons of the global model, TraceFL dynamically quantifies the significance of the global model’s neurons in a given prediction. It then selectively picks a slice of the most crucial neurons in the global model and maps them to the corresponding neurons in every participating client to determine each client’s contribution, ultimately localizing the responsible client. We evaluate TraceFL on six datasets, including two real-world medical imaging datasets and four neural networks, including advanced models such as GPT. TraceFL achieves 99% accuracy in localizing the responsible client in FL tasks spanning both image and text classification tasks. At a time when state-of-the-art ML debugging approaches are mostly domain-specific (e.g., image classification only), TraceFL is the first technique to enable highly accurate automated reasoning across a wide range of FL applications.
@@
Interpretability, Explainability, Debugging, Machine Learning, Federated Learning, Transformer, GPT
@@ I Introduction
Problem. In federated learning, the client(s) most responsible for a global model’s prediction are the ones trained on data that contains the predicted labels. This is analogous to finding influential training samples in classical machine learning [13]. However, the two domains, single model-based centralized ML and FL, are fundamentally orthogonal. Existing influence-based debugging approaches in ML [14, 15, 16, 17] and regular software [18, 19, 20], require transparent access to data including all data manipulation operations applied on the input data. When applied to FL, these approaches will require end-to-end monitoring of clients’ training (i.e., require access to clients’ data), which is prohibited in FL. More broadly, these ML influence and explainability-based debugging approaches target a single model in which the debugging is restricted to identifying the training data. In contrast, debugging in FL entails isolating a client’s model among many. This paper addresses the following debugging problem in FL: Given the global model inference on an input in FL, how can we identify the client(s) most responsible for the inference? Addressing it will have significant implications on faulty client identification, client incentivization, and the transparency and interpretability of FL.
@@
Challenges. Determining a client’s influence on the global model is challenging. Clients are randomly sampled in each round, each with unique data, contributing differently to the global model. Thus, the influence of a client on the global model is dynamic, non-uniform, and changes across rounds, making it difficult to link the global model’s behavior to a specific client. The FL protocol restricts access to client-side training, turning FL configuration into a nearly black-box setting. Additionally, clients’ models are collections of neuron weights that are individually uninterpretable. Static analysis of models’ weights to measure clients’ influence is ineffective because clients’ models are intrinsically different in terms of weights, making weight comparison futile. Furthermore, neural networks today comprise millions of neurons (e.g., GPT-3 has 175 billion parameters [21]) where tracing each neuron dynamically is infeasible.
@@
FL is increasingly used for domains other than vision using various neural network architectures, such as transformer and convolutional neural networks. Designing a generic FL debugging approach is a major challenge. For instance, transformers contain a self-attention mechanism that allows the model to focus on different parts of the input sequence. This mechanism is usually not seen in convolutional networks; instead, it uses a convolutional layer to detect the special patterns in the input data. Additionally, these architectures use different activation functions such as Rectified Linear Unit (ReLU) [22] and Gaussian Error Linear Unit (GELU) [23], introducing another source of complications.
@@
Our Contribution. We present the concepts of neuron provenance in FL, a fine-grained lineage-capturing mechanism that formulates the flow of information in the fusion algorithm from multiple clients’ models into a global FL model, ultimately influencing the predictions of the global model. Using neuron provenance, we determine the precise magnitude of contributions of participating clients towards the global model’s prediction. We materialize the idea of neuron provenance in TraceFL. TraceFL runs at the aggregator (i.e., central server) and requires no instrumentation on the client side.
@@
TraceFL is designed with the following insights. A global model consists of millions of neurons, making it infeasible to trace where how each neuron in a global model is computed. We observe that a dynamic subset of neurons activates in response to a given input, and not all neurons contribute equally to a prediction [24, 25]. Using this insight, TraceFL quantifies the contribution of these neurons in the global model’s prediction by computing the gradient of the neurons w.r.t. to the prediction. Such neuron-level gradients reveal the neuron’s output impact on the global model’s prediction and thus reduce the scope of important neurons. TraceFL then maps the global model’s important neurons to the corresponding neurons in each client’s model and computes the contribution of each client’s neuron to the corresponding global model’s neuron. At this stage, TraceFL computes the end-to-end neuron provenance of a global model prediction with the magnitude of the contribution of each client’s neurons. Finally, TraceFL aggregates the contributions of each client. The client with the highest contribution is deemed the most responsible for the given prediction.
@@
Evaluations. We demonstrate TraceFL’s effectiveness, generalizability, and robustness by evaluating its client localization accuracy on both image and language models under various commercial data distributions and differential privacy methods. We evaluate TraceFL on four state-of-the-art neural networks: ResNet [26], DenseNet [27], BERT [28], and GPT [29] and using six datasets including two real-world medical imaging datasets [30, 31, 32]. TraceFL achieves an average accuracy of 99% in localizing the responsible client across 30 unique FL settings, spanning both correct and misprediction scenarios. For fault localization in FL, TraceFL achieves an average accuracy of 99%, compared to 32% by the existing technique [33], demonstrating TraceFL’s effectiveness in real-world FL deployments. Additionally, we test TraceFL’s robustness against varying data distributions and differential privacy settings and find that TraceFL remains robust and effective, maintaining its 99% accuracy. We also vary the number of clients to up to 1000 and find that TraceFL is scalable and efficient in all settings. Overall, we evaluate TraceFL on 20,600 trained client models. These experiments exceed prior research’s evaluation complexity and fully represent commercial FL usage [34, 35, 36, 37, 38]. TraceFL is implemented in FLower FL [39] and compatible with GPU for parallel processing of neuron provenance for compute-intensive models (e.g., GPT).
@@
Novelty. TraceFL advances the state of FL debugging with the following core contributions:
@@
  • TraceFL introduces a unique concept of neuron-level provenance for FL applications to capture the dynamic contribution of each client, which helps rank clients based on the contribution to a given prediction. TraceFL efficiently tracks the contribution of clients in large models like GPT containing millions of parameters.
  • TraceFL achieves 99% localization accuracy in localizing the responsible client in FL. TraceFL localization remains intact during localizing a faulty client where existing baseline[33] achieves 32%.
  • TraceFL is the first approach that is equally effective on transformers and CNNs. Even the most sophisticated ML debugging approaches work on single model architecture and data domains, i.e., either convolutional neural networks (CNNs) or transformers.
  • TraceFL significantly advances the field of debugging and explainability in FL, addressing open challenges in FL [33, 40] and work with differential privacy and real-world data distributions among FL clients.
Source Code. TraceFL’s artifact is available at https://github.com/warisgill/TraceFL.
@@ II-A Federated Learning
It enables multiple clients (e.g., mobile devices, organizations) to train a shared model without sharing their data. This allows the model to be trained using distributed data, which can be useful in cases where data is distributed across multiple devices or organizations and cannot be easily collected and centralized. One algorithm of FL is Federated Averaging (FedAvg) [1], which uses the following equation to update the global model at each round of the training process:
@@ II-B Motivation
This section presents a real-world FL application that motivates the need for TraceFL in two common scenarios.
Incentivization. Suppose a developer deploys an FL system to diagnose colon diseases based on colon pathology images, as shown in Figure 1. In this FL system, ten hospitals, identified as H0 to H9, collaborate to train a global FL model. Each hospital trains its local model, which is then aggregated to form a global model. The classification stage in Figure 1 shows a scenario where the global model makes a correct prediction on new test colon pathology images (e.g., ‘Cancer-associated Stroma’ and ‘Mucus’). Since these are correct predictions, the FL developer aims to determine which hospital is most responsible for these correct predictions so that it can be encouraged to participate in future rounds. Since the training data is protected under the privacy of medical records and inaccessible to the developer, it is challenging to identify the responsible hospital by inspecting the raw model weights shared by the hospitals.
To address this issue, the developer decides to use TraceFL to identify the most responsible client behind the correct predictions. When enabled during the global model’s prediction, TraceFL localizes the hospital with H2 as the one responsible for the prediction of ‘Mucus’, and the hospital H6 as the one responsible for the prediction of ‘Cancer-associated Stroma’. More specifically, as shown on the right in Figure 1, TraceFL ranks hospitals (i.e., clients) based on their contributions to each prediction. The score associated with each hospital quantified how responsible a hospital is for that prediction. The training data distribution on the left shows that H2’s training data include data points labeled ‘Mucus’, whereas H6’s training data include data points labeled ‘Cancer-associated Stroma’.
Conversely, hospital H1’s contribution is 0.02 and 0.001 in the two predictions because it does not have any data points with the labels ‘Mucus’ or ‘Cancer-associated Stroma’, respectively. Detailed evaluations on two real-world medical imaging datasets are presented in Section V-A. Localizing the clients responsible for the prediction with TraceFL serves as a basis for designing advanced incentivization approaches.
Interpretability and Accountability. In addition to incentivization, TraceFL can effectively assist FL developers in localizing the hospital responsible for misprediction. Suppose the global model makes an incorrect prediction (e.g., predicts a cancer-associated stroma as mucus or vice versa). In such cases, FL developers can also use TraceFL to identify the root cause of the issue. Existing techniques  [41, 42, 43, 44, 45, 46, 47] are designed to interpret a single model and cannot be applied to the complex global models of FL, which are the fusion of many clients’ models. By utilizing TraceFL, FL developers gain valuable insights into the composition of the global model. TraceFL tracks the flow of information between hospitals and the global model at a fine-grained level. It can identify the most influential neurons of the global model for a given prediction and trace their origin to find the potential hospital(s) primarily responsible for incorrect predictions.
@@ III Challenges in Debugging FL
State-of-the-art neural networks, such as Transformers (BERT, GPT) and CNNs (ResNet, DenseNet), have varying structures, activation functions, and numbers of parameters. For example, GPT [29] is a 37-layer (12 block) Transformer architecture with Gaussian Error Linear Unit (GELU) [23] activation function and 117 million parameters.DenseNet [27] has 121 layers, 8 million parameters, and uses the Rectified Linear Unit (ReLU) [22] activation function. For instance, in a single FL round with 10 clients using GPT, 10$\times$117 million parameters are fused to form a new global model. The fusion algorithm or the global model does not inherently provide any information about individual clients’ contributions. Thus, tracking a client’s contribution among millions of parameters is a significantly challenging task.
Moreover, the data distribution across FL training rounds is non-identical; clients rarely share the same data point. Not all clients participate in every round, and some clients may contain only a few data points to train their local model. The class label distribution also varies across clients. Such variability causes more hurdles in precisely reasoning about a global model’s behavior. Simply tracking the static weights of the global model is inadequate, as different sets of neurons are activated on different inputs. Inspecting individual clients’ models does not help understand the client’s contribution to the global model prediction, as it will not capture the cumulative behavior of the global model.
@@
8 for each activated neuron in global_model do
@@ IV Design of TraceFL
TraceFL addresses the aforementioned challenges using neuron provenance. At a high level, TraceFL dynamically tracks the lineage of the global model at the neuron level and identifies the most influential clients against a given prediction by the global model ($W_{global}^{t+1}$) on an input. Enabling provenance at the neuron level solves the complexities of different neural networks architectural design (e.g., number of layers and parameters, different activation functions) and enable it to work in cross domains such as image and text classification tasks.
TraceFL comprises the following steps: First, TraceFL identifies the activated influential neurons in the global model for a given test input. Next, TraceFL traces the contribution of each client to all the activated neurons. Finally, TraceFL computes the total contribution of each client to the global model. These steps collectively construct a comprehensive end-to-end provenance graph, which is used to debug the contributions of clients in the given prediction of the global model. Algorithm 1 outlines the design of TraceFL.
@@ IV-A Determining Influential Neurons
TraceFL first aims to identify neurons that actively participate in an FL global model’s prediction. Traditional data provenance approaches must trace the participation of all input data records in the operation for completeness, eventually mapping them to individual outputs of the operation. However, tracking the provenance of all neurons is wasteful because not all neurons participate equally in a model’s prediction. Therefore, tracking the behavior of all neurons in a model may lead to over-approximation (i.e., more than expected clients are classified as contributors) when provenance is used to identify the contributing clients.
@@
The behavior of a neural network on a given input is determined by the set of activated neurons in the network, and different sets of neurons are activated on different inputs. We leverage this insight and apply TraceFL’s neuron provenance to dynamically quantify the influence of global model neurons on each prediction for the given input. This reduce the likelihood of over-approximation by excluding the neurons that may distort the outcome when the lineage of a specific neuron is used to localize the influential client.
@@
Mathematically, the output of a neuron is $z=\sigma(\mathbf{w}\cdot\mathbf{z})$, where $\mathbf{w}$ is the set of weights of the neuron, $\mathbf{z}$ is the input to a neuron, and $\sigma$ is the activation function. One of the commonly used activation functions ($\sigma$) is ReLU (Rectified Linear Unit) [22]. The output of $\sigma$ is called the activation or output of the neuron. A neuron with ReLU function is consider active if $z>0$. Note that the output of a neuron ($z$) is part of the input to the neurons of the next layer. Next, TraceFL computes the activation of each neuron in the network. Suppose that $n_{j}$ represents the $j$-th neuron in a neural network and the set of all the outputs (i.e., activations) of all the neurons in a neural network can be represented as $\{z_{n_{1}},z_{n_{2}},...,z_{n_{j}}\}$, which captures the complete dynamic behavior of the network on a given input $\mathbf{x}$. Note that for the first layer, the input $\mathbf{z}$ to neurons will be the $\mathbf{x}$ i.e., $\mathbf{z}=\mathbf{x}$ for the first layer of the neural network.
@@
After computing the global model neurons activations, TraceFL’s goal is to find their measurable contribution towards the global model’s prediction ($y$) on an input. In the output ($y$) of the global model, not all the activated neurons carry equal importance. For instance, neurons in the last layers learn better and more rigorous features than neurons in the initial layers of the network [24]. Since TraceFL aims to localize the client that contributed the most towards a prediction, assigning equal importance to all activated neurons will again cause over-approximation or even wrong client localization. To enable precise and accurate provenance, we must measure the individual influence of a neuron on the final prediction.

### v4

@@ Abstract
In Federated Learning, clients train models on local data and send updates to a central server, which aggregates them into a global model using a fusion algorithm. This collaborative yet privacy-preserving training comes at a cost. FL developers face significant challenges in attributing global model predictions to specific clients. Localizing responsible clients is a crucial step towards (a) excluding clients primarily responsible for incorrect predictions and (b) encouraging clients who contributed high-quality models to continue participating in the future. Existing ML debugging approaches are inherently inapplicable as they are designed for single-model, centralized training.
We introduce TraceFL, a fine-grained neuron provenance capturing mechanism that identifies clients responsible for a global model’s prediction by tracking the flow of information from individual clients to the global model. Since inference on different inputs activates a different set of neurons of the global model, TraceFL dynamically quantifies the significance of the global model’s neurons in a given prediction, identifying the most crucial neurons in the global model. It then maps them to the corresponding neurons in every participating client to determine each client’s contribution, ultimately localizing the responsible client. We evaluate TraceFL on six datasets, including two real-world medical imaging datasets and four neural networks, including advanced models such as GPT. TraceFL achieves 99% accuracy in localizing the responsible client in FL tasks spanning both image and text classification tasks. At a time when state-of-the-art ML debugging approaches are mostly domain-specific (e.g., image classification only), TraceFL is the first technique to enable highly accurate automated reasoning across a wide range of FL applications.
@@
Interpretability, Explainability, Debugging, Machine Learning, Federated Learning, Transformer
@@ I Introduction
Problem. In federated learning, the client(s) most responsible for a global model’s prediction are the ones trained on data that contains the predicted labels. This is analogous to finding influential training samples in classical machine learning [13]. However, the two domains, single model-based centralized ML and FL, are fundamentally different. Existing influence-based debugging approaches in ML [14, 15, 16, 17] and regular software [18, 19, 20] require transparent access to data including all data manipulation operations applied on the input data. When applied to FL, these approaches will require end-to-end monitoring of clients’ training (i.e., require access to clients’ data), which is prohibited in FL. More broadly, ML influence and interpretability-based debugging approaches target a single model in which the debugging is restricted to identifying the training data. In contrast, debugging in FL entails isolating a client’s model among many. This paper addresses the following debugging problem in FL: Given the global model inference on an input in FL, how can we identify the client(s) most responsible for the inference?
@@
Challenges. Determining a client’s influence on the global model is challenging. Clients are randomly sampled in each round, each possessing unique data and contributing differently to the global model. Thus, the influence of a client on the global model is dynamic, non-uniform, and changes across rounds, making it difficult to link the global model’s behavior to a specific client. The FL protocol restricts access to client-side training, turning FL configuration into a nearly black-box setting. Additionally, clients’ models are collections of neuron weights that are individually uninterpretable. Static analysis of models’ weights to measure clients’ influence is ineffective because clients’ models are intrinsically different in terms of weights. Furthermore, neural networks today comprise millions of neurons (e.g., GPT-3 has 175 billion parameters [21]). Considering all neurons equally, in such cases, would lead to imprecise and incorrect debugging.
@@
FL is increasingly used for domains other than vision using various neural networks, such as transformer and convolutional neural networks (CNNs). Designing a generic FL debugging approach is a major challenge. For instance, transformers contain a self-attention mechanism that allows the model to focus on different parts of the input sequence. This mechanism is usually not seen in CNNs; instead, it uses a convolutional layer to detect the special patterns in the input data. Additionally, these architectures use different activation functions such as Rectified Linear Unit (ReLU) [22] and Gaussian Error Linear Unit (GELU) [23], introducing another source of complications.
@@
Our Contribution. We present the concepts of neuron provenance, a fine-grained lineage-capturing mechanism that formulates the flow of information in the fusion algorithm from multiple clients’ models into a global FL model, ultimately influencing the predictions of the global model. Using neuron provenance, we determine the precise magnitude of contributions of participating clients towards the global model’s prediction. We materialize the idea of neuron provenance in TraceFL, which runs at the aggregator (i.e., central server) and requires no instrumentation on the client side.
@@
TraceFL is designed with the following insights. Since a global model consists of millions of neurons, we observe that a dynamic subset of neurons activates in response to a given input, and not all neurons contribute equally to a prediction [24, 25]. Using this insight, TraceFL quantifies the contribution of these neurons in the global model’s prediction by computing the gradient of the neurons w.r.t. to the prediction. Such neuron-level gradients reveal the neuron’s output impact on the global model’s prediction and thus reduce the scope of important neurons.
TraceFL then maps the global model’s important neurons to the corresponding neurons in each client’s model and computes the contribution of each client’s neuron to the corresponding global model’s neuron. At this stage, TraceFL computes the end-to-end neuron provenance of the global model prediction with the magnitude of the contribution of each client’s neurons. Finally, TraceFL aggregates the contributions of each client. The client with the highest contribution is deemed the most responsible for the given prediction.
@@
Evaluations. We demonstrate TraceFL’s effectiveness, generalizability, and robustness by evaluating its client localization accuracy on both image and language models under various commercial data distributions and differential privacy methods. We evaluate TraceFL on four state-of-the-art neural networks: ResNet [26], DenseNet [27], BERT [28], and GPT [29] and using six datasets including two real-world medical imaging datasets [30, 31, 32]. TraceFL achieves an average accuracy of 99% in localizing the responsible client across 30 unique FL settings, spanning both correct and misprediction scenarios. For fault localization in FL, TraceFL achieves an average accuracy of 99%, compared to 32% by the existing technique [33], demonstrating TraceFL’s effectiveness in real-world FL deployments. Additionally, we test TraceFL’s robustness against varying data distributions and differential privacy settings and find that TraceFL remains robust and effective. We also vary the number of clients, increasing it up to 1000, and find that TraceFL is both scalable and efficient. Overall, we evaluate TraceFL on 20,600 trained client models. These experiments exceed prior research’s evaluation complexity and fully represent commercial FL usage [34, 35, 36, 37, 38]. TraceFL is implemented in Flower FL [39] and compatible with GPU for parallel processing of neuron provenance for compute-intensive models (e.g., GPT).
@@
TraceFL advances the state of FL debugging with the following core contributions:
@@
  • TraceFL introduces a unique concept of neuron provenance for FL applications to capture the dynamic contribution of each client, which helps rank clients based on the contribution to a given prediction. TraceFL efficiently tracks the contribution of clients in large models like GPT containing millions of parameters.
  • TraceFL achieves 99% localization accuracy in localizing the responsible client in FL. TraceFL’s localization accuracy remains high during localizing a faulty client where existing baseline[33] achieves 32%.
  • TraceFL is the first approach that is equally effective on transformers and CNNs. Even the most sophisticated ML debugging approaches work on single model architecture and data domains, i.e., either CNNs or transformers.
  • TraceFL significantly advances the field of debugging and interpretability in FL, addressing open challenges in FL [33, 40] and work with differential privacy and real-world data distributions among FL clients.
Source Code. TraceFL’s artifact is available at https://github.com/SEED-VT/TraceFL.
@@ II-A Federated Learning
Federated Learning enables multiple clients (e.g., mobile devices, organizations) to train a shared model without sharing their data. This allows the model to be trained using distributed data, which can be useful in cases where data is distributed across multiple devices or organizations and cannot be easily collected and centralized. One algorithm of FL is Federated Averaging (FedAvg) [1], which uses the following equation to update the global model at each round of the training process:
@@
FL has variations such as Vertical FL [41] and Personalized FL [42]. TraceFL primarily focuses on horizontal FL [1], similar to previous work on fault localization in FL [33, 10]. A typical FL setup involves a few to thousands of clients, such as mobile devices, healthcare institutions, or enterprises. FL clients exhibit diversity in data distribution and computational resources. Data is often non-independently and identically distributed (Non-IID), with varying sizes across clients, such as hospitals specializing in different medical conditions (Figure 1). All participating clients use the same neural network architecture, ensuring compatibility during model aggregation. Additionally, clients have heterogeneous hardware and network capabilities (e.g., smartphones to powerful servers), impacting their participation and training consistency [43].
@@ II-B Motivation
Suppose a developer deploys an FL system to diagnose colon diseases based on colon pathology images, as shown in Figure 1. In this FL system, ten hospitals, identified as H0 to H9, collaborate to train a global FL model. Each hospital trains its local model, which is then aggregated to form a global model. The classification stage in Figure 1 shows a scenario where the global model makes a correct prediction on new test colon pathology images (e.g., ‘Cancer-associated Stroma’ and ‘Mucus’). Since these are correct predictions, the FL developer aims to determine which hospital is most responsible for these correct predictions so that it can be encouraged to participate in future rounds. Since the training data is protected under the privacy of medical records and inaccessible to the developer, it is challenging to identify the responsible hospital by inspecting the raw model weights shared by the hospitals.
To address this issue, the developer decides to use TraceFL to identify the most responsible client behind the correct predictions. When enabled during the global model’s prediction, TraceFL localizes the hospital with H2 as the one responsible for the prediction of ‘Mucus’, and the hospital H6 as the one responsible for the prediction of ‘Cancer-associated Stroma’. More specifically, as shown on the right in Figure 1, TraceFL ranks hospitals (i.e., clients) based on their contributions to each prediction. The score associated with each hospital quantifies how responsible a hospital is for that prediction. The training data distribution on the left shows that H2’s training data include data points labeled ‘Mucus’, whereas H6’s training data include data points labeled ‘Cancer-associated Stroma’.
Conversely, hospital H1’s contribution is 0.02 and 0.01 in the two predictions because it does not have any data points with the labels ‘Mucus’ or ‘Cancer-associated Stroma’, respectively. Detailed evaluations on two real-world medical imaging datasets are presented in Section V-A. Localizing the clients responsible for the prediction with TraceFL serves as a basis for designing advanced incentivization approaches.
@@ III Challenges in Debugging FL
State-of-the-art neural networks, such as Transformers (BERT, GPT) and CNNs (ResNet, DenseNet), have varying structures, activation functions, and numbers of parameters. For example, GPT [29] is a 37-layer (12 block) Transformer architecture with GELU [23] activation function and 117 million parameters. DenseNet [27] has 121 layers, 8 million parameters, and uses the ReLU [22] activation function. The fusion algorithm or the global model does not inherently provide any information about individual clients’ contributions. Thus, tracking a client’s contribution among millions of parameters is a significantly challenging task.
Moreover, the data distribution across FL training rounds is non-identical; clients rarely have the same data point. Not all clients participate in every round, and some clients may contain only a few data points to train their local model. The class label distribution also varies across clients. Such variability causes more hurdles in precisely reasoning about a global model’s behavior. Simply tracking the static weights of the global model is inadequate, as different sets of neurons are activated on different inputs, and not all neurons have equal importance. Inspecting individual clients’ models does not help understand the client’s contribution to the global model prediction, as it will not capture the cumulative behavior of the global model.
@@
8 for each neuron in global_model do
@@ IV Design of TraceFL
TraceFL addresses the aforementioned challenges using neuron provenance. At a high level, TraceFL dynamically tracks the lineage of the global model at the neuron level and identifies the most influential clients against a given prediction by the global model ($W_{global}^{t+1}$) on an input. Enabling provenance at the neuron level solves the complexities of different neural networks architectural design (e.g., number of layers and parameters, different activation functions) and enables TraceFL to work in cross domains such as image and text classification tasks.
In essence, TraceFL first identifies the influential neurons by jointly analyzing neuron activations, the layers the neurons are in, and gradients with respect to individual neurons in the global model for a given test input. Next, TraceFL precisely quantifies the individual contribution of each corresponding neuron from every participating client to the neurons in the global model. Finally, TraceFL computes the total contribution of each client to the global model. These steps collectively construct a comprehensive end-to-end provenance graph, which is used to debug the contributions of clients in the given prediction of the global model. Algorithm 1 outlines the design of TraceFL.
@@ IV-A Determining Influential Neurons
TraceFL first aims to identify neurons that actively participate in an FL global model’s prediction. Traditional data provenance approaches must trace the participation of all input data records in the operation for completeness, eventually mapping them to individual outputs of the operation. However, tracking the provenance of all neurons with equal importance is wasteful because not all neurons participate equally in a model’s prediction. Therefore, tracking the behavior of neurons with the same importance in a model may lead to over-approximation (i.e., more than expected clients are classified as contributors) when provenance is used to identify the contributing clients.
@@
The behavior of a neural network on a given input is determined by the set of activated neurons in the network, and different sets of neurons are activated on different inputs. We leverage this insight and apply TraceFL’s neuron provenance to dynamically quantify the influence of global model neurons on each prediction for the given input. This reduces the likelihood of over-approximation by minimizing the contribution of neurons that may distort the outcome when the lineage of a specific neuron is used to localize the influential client.
@@
Mathematically, the output of a neuron is $z=\sigma(\mathbf{w}\cdot\mathbf{z})$, where $\mathbf{w}$ is the set of weights of the neuron, $\mathbf{z}$ is the input to a neuron, and $\sigma$ is the activation function. One of the commonly used activation functions ($\sigma$) is ReLU  [22]. The output of $\sigma$ is called the activation or output of the neuron. A neuron with ReLU function is considered active if $z>0$. Note that the output of a neuron ($z$) is part of the input to the neurons of the next layer. Next, TraceFL computes the activation of each neuron in the network. Suppose that $n_{j}$ represents the $j$-th neuron in a neural network and the set of all the outputs (i.e., activations) of all the neurons in a neural network can be represented as $\{z_{n_{1}},z_{n_{2}},...,z_{n_{j}}\}$, which captures the complete dynamic behavior of the network on a given test input $\mathbf{x}$. Note that for the first layer, the input $\mathbf{z}$ to neurons will be the input $\mathbf{x}$ to the model i.e., $\mathbf{z}=\mathbf{x}$ for the first layer of the neural network.
@@
After computing the global model neurons activations, TraceFL’s goal is to find their measurable contribution towards the global model’s prediction ($y$) on an input. In the output ($y$) of the global model, not all the neurons carry equal importance. For instance, neurons in the last layers learn better and more rigorous features than neurons in the initial layers of the network [24]. Since TraceFL aims to localize the client that contributed the most towards a prediction, assigning equal importance to all neurons will again cause over-approximation or even wrong client localization. To enable precise and accurate provenance, we must measure the individual influence of a neuron on the final prediction.